OpenStudy (anonymous):

CO(g)+NH3(g)⇌HCONH2(g), Kc=0.800 If a reaction vessel initially contains only CO and NH3 at concentrations of 1.00 M and 2.00 M, respectively, what will the concentration of HCONH2 be at equilibrium?

OpenStudy (abmon98):

[Products]^n/[Reactants]^n=Kc Initial Concentration: 1.00 2.00 - At equilbrium: 1-x, 2-x, x x/(1-x)(2-x)=0.800 x/1-x-2x+x^2=0.800 x/1-3x+x^2=0.800 0.800(1-3x+x^2)=x